# ReactEMG: Zero-Shot, Low-Latency Intent Detection via sEMG

### v3

@@ 1 Introduction
    Limited lookahead for low latency: consider an intent detection model that, in order to classify the intent at time $t$, uses EMG data collected up to time $t+\ell$ (Fig. 1 Middle). Using $\ell>0$ can be beneficial for the model, providing additional data as the new gesture unfolds. However, since the prediction must now lag behind ground truth intent change by a time interval of at least length $\ell$, using large values for $\ell$ introduces latency in the response. We thus focus on methods where the value of $\ell$, dubbed here "lookahead" time, is low enough to ensure responsiveness.
@@
    Stability: while the ideal model must recognize successive intent changes, it must also deal appropriately with long "maintenance windows" where a single intent is maintained for longer time periods. During such windows, inadvertent "flicker" in the detected intent can have negative consequences if it produces corresponding flicker in the device being controlled. We thus focus on methods that distinguish between transient activity at a gesture’s onset and sustained activation throughout its maintenance (Fig. 1 Right), and do not predict false transitions during the latter.
These desired characteristics are partially captured by the traditional metric of per-time-step classification accuracy, dubbed here "raw accuracy". While good performance on raw accuracy is important, it can also hide situations where a transition to a new gesture is detected with delay, or a maintenance window exhibits unwanted flicker. To capture such cases, we use a new metric dubbed "transition detection accuracy". According to this metric, a transition is considered to be successfully detected only if the model output correctly transitions between intents close enough to the ground truth change, and exhibits no instability either before or after the transition.
@@ 2 Related work
Beyond core intent detection tasks, several large-scale sEMG benchmarks have been released to investigate broader challenges in muscle-driven interfaces. Sivakumar et al. introduce a dataset of wrist sEMG from 108 users paired with keystroke transcripts for speech-recognition–style sequence modeling, but observe a sharp performance drop when generalizing to unseen users without adaptation [33]. Salter et al. introduces a dataset of 16-channel HD-sEMG synchronized to 26-camera motion capture across 193 users and 29 gesture stages [34]. Their pipeline excels at reconstructing continuous joint kinematics, though it does not emphasize higher-level semantic intent recognition, which need to accommodate the natural variations in how different individuals perform the same gesture. Meanwhile, Yang et al. aggregate nine public gesture corpora to benchmark out-of-distribution classification via leave-one-subject-out and few-shot adaptation, but do not address streaming, low-latency decoding or rapid calibration-free deployment [35].
@@
Most existing methods focus on gesture recognition with retrospective classification, leading to delays and a lack of real-time performance. Our method, however, provides continuous intent detection with low-latency predictions at every timestep. We introduce a masked modeling strategy that aligns EMG signals with user intent, enabling faster and more accurate predictions. Unlike existing approaches that require extensive calibration for each user, our model generalizes to new subjects without additional training. Additionally, while other methods suffer from instability or "flickering" during sustained gestures, our approach maintains stable intent detection over long periods. These differences make our method more efficient and reliable for real-world applications.
@@ 3.3 Online Inference
Lookahead for smoother predictions. Suppose we want to predict the intent at timestep $t$. To do this, we introduce a small lookahead parameter $\ell$, which allows us to collect logits for timesteps from $t$ to $t+\ell$ across subsequent windows. We then aggregate the model outputs from all windows that cover the interval $[t,\,t+\ell]$ by averaging their logits. Finally, we obtain the predicted class label for timestep $t$ by applying argmax to this merged distribution.
@@ 4.1 Metrics
Figure 3: Illustration of the Transition Accuracy evaluation protocol. For each ground-truth change from class $y_{\mathrm{old}}$ to $y_{\mathrm{new}}$, we define a reaction buffer (blue region) centered on the ground truth transition. The model must predict $y_{\mathrm{new}}$ at least once within this buffer and must not predict any other class. The subsequent maintenance period (red region) extends from the end of the reaction buffer to the start of the next reaction buffer; during this interval, the model’s output must remain stably equal to $y_{\mathrm{new}}$. A transition is scored as correct only if both reaction buffer and maintenance period conditions are satisfied.
@@
Let the ground-truth label at time $t$ be $y_{t}$. We define a "transition" as a segment of data that begins when the label transitions from a previous class $y_{\mathrm{old}}$ to a new class $y_{\mathrm{new}}$, and ends at the next class switch. Each transition is divided into two disjoint intervals: (1) Reaction buffer: A short time window (typically 1 s in our experiments) centered on the ground truth intent transition timestep. The model must predict transition from $y_{\mathrm{old}}$ to $y_{\mathrm{new}}$ at least once within this buffer, and should not contain any other prediction classes. We use a buffer window in our transition accuracy metric to account for the fact that the exact timestamp of a transition is very difficult to identify even in ground truth data (since the time elapsed between an experimenter issuing an instruction and the subject executing the corresponding gesture can vary, and offline, manual labeling is subjective and impractical for large datasets). However, since we want models to be responsive in real-life use, we keep the buffer window short. (2) Maintenance period: The maintenance period spans the remainder of the gesture until the onset of the next reaction buffer. After predicting $y_{\mathrm{new}}$, the model must consistently output the same label for the entire duration of this interval. Any incorrect prediction within this interval immediately invalidates the entire transition. Predictions on a transition is considered correct if and only if both these conditions are met. By penalizing both delayed onset detection and instability during the maintenance period, this metric provides a stringent and realistic measure of performance in real-world use cases.
@@ 7 Acknowledgments
Appendix A Online Inference
@@
Figure 4 illustrates our method as used for online inference. In order to label the intent at time $t$, we take all the overlapping windows of size equal to 600 timesteps = $3s$ and whose final sample falls inside the look-ahead horizon $[t,t+\ell]$ (yellow region, $\ell=50$ timesteps = 250 ms), and feed them through our model. Our model then produces one logit for every timestep of each of these windows. We then average all the logits for timesteps falling within the lookahead horizon, and apply an $\arg\max$ to the result in order to obtain a single class label for time step $t$. To avoid jitter, the resulting label from time $t$ is then maintained for a hold period of $s=20$ timesteps or 100 ms, (blue region), after which the online inference process is performed again for time $t+100ms$.
@@
We note that our approach inherently introduces two types of latency in the response. First, the label for time $t$ is issued after ingesting data up to $t+250ms$, thus introducing up to 250ms delay in the response to a change in intent. However, this delay is smaller than the inherent delay in human reaction time to visual or verbal cues (which we discuss further in Section F). Second, the hold period means that our model will not react to intent changes happening at a rate faster than 10 Hz. However, we assume it to be highly unlikely that a human subject will modulate their hand pose or intent at such a rate.
Appendix B Dataset Description
@@
Figure 5 compares representative predictions from two models: one trained only on the public datasets and another pretrained on the public datasets and fine-tuned on our data. We evaluate both models on two tasks from our dataset: unsupported static hand gestures and functional grasp of a soup can. As shown, while both models initially identify the correct intent, the pretrain-only model quickly reverts to a "rest" state. In contrast, the fine-tuned model maintains the correct gesture label throughout the task, demonstrating improved temporal consistency and real-world applicability.
Appendix C Model Hyperparameters and Training Configurations
@@
Appendix D Baseline Implementations
@@
Betthauser et al. [17] proposes ED-TCN, an encoder–decoder temporal convolutional network that relies solely on 1D convolutions. To establish a fair basis of comparison, all baseline models use 600 window size, except ED-TCN. To align with the original paper’s hyperparameters, we increase the ED-TCN context length from 600 to 625 samples. The model processes these 625-sample “outer windows” with a stride of 30. Within each outer window, we then slide a 150-sample “inner window” (stride 25), producing 20 sub-windows (as opposed to 19 if the length remained 600). For each sub-window, we compute the mean absolute value (MAV) across all eight EMG channels, resulting in a 20×8 tensor that serves as input to ED-TCN. We maintain the original two-layer encoder structure, where each layer consists of a one-dimensional convolution, ReLU activation, and 2× max pooling. The decoder mirrors this architecture by applying 2× upsampling and the same convolution+ReLU sequence, concluding with a 1×1 convolution that outputs class probabilities for each time step. To accommodate the 625-sample windows, we scale down the convolution kernel size proportionally, using a kernel size of 9. This ensures the final temporal length is exactly 20 after pooling and upsampling. During training, we compute cross-entropy loss at each sub-window and backpropagate through every time step (where each time step corresponds to the MAV values of a sub-window), and use the AdamW optimizer (learning rate = 1e-4) to mitigate overfitting. This slight adjustment—using 625-sample windows instead of 600 and a kernel size of 9 instead of 25—aligns the number of sub-windows with ED-TCN’s down/up-sampling factors, ensuring a one-to-one mapping between sub-window features and labels without additional padding or cropping.
@@
Appendix E Context Window Size Ablation
Window Size. Overall window size controls the amount of context that the model sees when making a prediction. We vary the window size from 100 timesteps (0.5s) to 600 timesteps (3s) to assess how temporal context affects intent detection. Table 5 shows results of our model using three different window sizes on the EMG-EPN-612 6-class dataset, with the evaluation configured to use lookahead 0 so that we can directly compare raw performance without any smoothing effects. Smaller windows reduce predictive stability, likely because the model lacks sufficient information to reliably distinguish transitory muscle signals. Consequently, we select a 600-timestep window size in our final model.
@@ Appendix E Context Window Size Ablation
Appendix F Reaction Time Discussion
@@
Ground truth intent labels are typically provided by a human experimenter, who cues the subject to execute a given command, while simultaneously marking the new command as ground truth intent. However, human reaction times introduce a baseline delay in muscle activation relative to visually or verbally cued labels. Simple reaction times to visual stimuli commonly range from about 200–250 ms, though they can be shorter in younger adults or under optimal conditions, and can approach 300 ms if the individual is older, fatigued, or otherwise slower to respond [Woods2015, Jain2015]. Auditory cues tend to elicit slightly faster reaction times (around 150–200 ms) because neural processing for sound is somewhat quicker than for visual signals [Arafat2022Comparison, Jain2015]. Furthermore, participants may anticipate cues, activating muscles before or after the ground-truth label shift, which also misaligns the recorded EMG signal with the labeled transition. Manually labeling EMG datasets offline is imprecise and subjective, particularly when transitioning between non-rest states. Consequently, training and validation data can contain a baseline offset of 200–300 ms between the true transition and the labeled one. To account for this, we report results on our transition accuracy metric using a default reaction buffer size of 100 timesteps $=500ms$ on each side of the ground truth labeled transition, comparable to the natural variability in reaction times and labeling inconsistencies. We also note that the same buffer size is used for all compared models, both ours and baselines.
@@
We also note that our model adds only negligible computational delay, beyond the lookahead window discussed in Sec. A, as it requires no additional data post processing. On an Nvidia RTX 4060 laptop GPU, the forward inference time is approximately 3 ms, which we consider negligible compared to the physiological and labeling-induced delays discussed above.
@@
Appendix G Random Seed Experiments
In Table 7, we show raw accuracy and transition accuracy results averaged over five different random seeds, along with standard deviation, for both our method and all baselines, evaluated on the EPN 6-class dataset, with no lookahead. We notice that all models appear largely insensitive to the choice of random seed.

### v4

@@ 1 Introduction
    Limited lookahead for low latency: consider an intent detection model that, in order to classify the intent at time $t$, uses EMG data collected up to time $t+\ell$ (Fig. 1 Middle). Using $\ell>0$ can be beneficial for the model, providing additional data as the new gesture unfolds. However, since the prediction must now lag behind ground truth intent change by a time interval of at least length $\ell$, using large values for $\ell$ introduces latency in the response. We thus focus on methods where the value of $\ell$, dubbed here “lookahead” time, is low enough to ensure responsiveness.
@@
    Stability: while the ideal model must recognize successive intent changes, it must also deal appropriately with long “maintenance windows” where a single intent is maintained for longer time periods. During such windows, inadvertent “flicker” in the detected intent can have negative consequences if it produces corresponding flicker in the device being controlled. We thus focus on methods that distinguish between transient activity at a gesture’s onset and sustained activation throughout its maintenance (Fig. 1 Right), and do not predict false transitions during the latter.
These desired characteristics are partially captured by the traditional metric of per-time-step classification accuracy, dubbed here “raw accuracy”. While good performance on raw accuracy is important, it can also hide situations where a transition to a new gesture is detected with delay, or a maintenance window exhibits unwanted flicker. To capture such cases, we use a new metric dubbed “transition detection accuracy”. According to this metric, a transition is considered to be successfully detected only if the model output correctly transitions between intents close enough to the ground truth change, and exhibits no instability either before or after the transition.
@@ 2 Related work
Beyond core intent detection tasks, several large-scale sEMG benchmarks have been released to investigate broader challenges in muscle-driven interfaces. Sivakumar et al. introduce a dataset of wrist sEMG from 108 users paired with keystroke transcripts for speech-recognition–style sequence modeling, but observe a sharp performance drop when generalizing to unseen users without adaptation [33]. Salter et al. introduces a dataset of 16-channel HD‐sEMG synchronized to 26-camera motion capture across 193 users and 29 gesture stages [34]. Their pipeline excels at reconstructing continuous joint kinematics, though it does not emphasize higher-level semantic intent recognition, which need to accommodate the natural variations in how different individuals perform the same gesture. Meanwhile, Yang et al. aggregate nine public gesture corpora to benchmark out‐of‐distribution classification via leave‐one‐subject‐out and few‐shot adaptation, but do not address streaming, low‐latency decoding or rapid calibration‐free deployment [35].
@@
Most existing methods focus on gesture recognition with retrospective classification, leading to delays and a lack of real-time performance. Our method, however, provides continuous intent detection with low-latency predictions at every timestep. We introduce a masked modeling strategy that aligns EMG signals with user intent, enabling faster and more accurate predictions. Unlike existing approaches that require extensive calibration for each user, our model generalizes to new subjects without additional training. Additionally, while other methods suffer from instability or “flickering” during sustained gestures, our approach maintains stable intent detection over long periods. These differences make our method more efficient and reliable for real-world applications.
@@ 3.3 Online Inference
Lookahead for smoother predictions. Suppose we want to predict the intent at timestep $t$. To do this, we introduce a small lookahead parameter $\ell$, which allows us to collect logits for timesteps from $t$ to $t+\ell$ across subsequent windows. We then aggregate the model outputs from all windows that cover the interval $t,\,t+\ell]$ by averaging their logits. Finally, we obtain the predicted class label for timestep $t$ by applying argmax to this merged distribution.
@@ 4.1 Metrics
Figure 3: Illustration of the Transition Accuracy evaluation protocol. For each ground‐truth change from class $y_{\mathrm{old}}$ to $y_{\mathrm{new}}$, we define a reaction buffer (blue region) centered on the ground truth transition. The model must predict $y_{\mathrm{new}}$ at least once within this buffer and must not predict any other class. The subsequent maintenance period (red region) extends from the end of the reaction buffer to the start of the next reaction buffer; during this interval, the model’s output must remain stably equal to $y_{\mathrm{new}}$. A transition is scored as correct only if both reaction buffer and maintenance period conditions are satisfied.
@@
Let the ground-truth label at time $t$ be $y_{t}$. We define a “transition” as a segment of data that begins when the label transitions from a previous class $y_{\mathrm{old}}$ to a new class $y_{\mathrm{new}}$, and ends at the next class switch. Each transition is divided into two disjoint intervals: (1) Reaction buffer: A short time window (typically 1 s in our experiments) centered on the ground truth intent transition timestep. The model must predict transition from $y_{\mathrm{old}}$ to $y_{\mathrm{new}}$ at least once within this buffer, and should not contain any other prediction classes. We use a buffer window in our transition accuracy metric to account for the fact that the exact timestamp of a transition is very difficult to identify even in ground truth data (since the time elapsed between an experimenter issuing an instruction and the subject executing the corresponding gesture can vary, and offline, manual labeling is subjective and impractical for large datasets). However, since we want models to be responsive in real-life use, we keep the buffer window short. (2) Maintenance period: The maintenance period spans the remainder of the gesture until the onset of the next reaction buffer. After predicting $y_{\mathrm{new}}$, the model must consistently output the same label for the entire duration of this interval. Any incorrect prediction within this interval immediately invalidates the entire transition. Predictions on a transition is considered correct if and only if both these conditions are met. By penalizing both delayed onset detection and instability during the maintenance period, this metric provides a stringent and realistic measure of performance in real-world use cases.
@@ 7 Acknowledgments
Appendix A Masking Logic
A.1 Notation & Data Partitions
  • EMG window: $x\in{R}^{T\times C}$ with $T=600$ time‑steps and $C=8$ channels.
  • Action labels: $y\in\{0,\dots,K-1\}^{T}$ with $K$ classes.
  • Mask sets:
    $\mathcal{M}_{E}\subseteq\{1,\dots,T\}\times\{1,\dots,C\},\qquad\mathcal{M}_{A}\subseteq\{1,\dots,T\}.$
  • Dataset split: $\mathcal{D}=\mathcal{D}_{\mathrm{labeled}}\cup\mathcal{D}_{\mathrm{unlabeled}}.$
Task Objective Masks created Active loss terms
0 Predict action $\mathcal{M}_{A}\!\neq\!\varnothing$, $\mathcal{M}_{E}=\varnothing$ CE on $\mathcal{M}_{A}$
1 Predict EMG $\mathcal{M}_{E}\!\neq\!\varnothing$, $\mathcal{M}_{A}=\varnothing$ MSE on $\mathcal{M}_{E}$
2 Predict action and EMG both non‑empty CE + MSE
3 Self‑sup. EMG $\mathcal{M}_{E}\!\neq\!\varnothing$, $\mathcal{M}_{A}=\varnothing$ MSE only
Each labeled sample is replicated once per supervised task $t\in\{0,1,2\}$; every unlabeled sample participates only in the self‑supervised task ($t=3$).
A.2 Mask Type Mixture
For every sample, we draw
$\Pr(\text{Poisson})=0.99,\Pr(\text{End})=0.01,$
except that task 3 always uses Poisson (Sec. A.6). For the chosen mask type we sample
$p_{t}\sim\begin{cases}\mathrm{U}(0.30,0.75),&\text{Poisson},\;t\in\{0,1,2\},\\[4.0pt] \mathrm{U}(0.30,0.60),&\text{End},\;t\in\{0,1,2\},\\[4.0pt] \mathrm{U}(0.20,0.65),&\text{Poisson},\;t=3,\end{cases}\quad\Longrightarrow\quad N_{t}=\lfloor p_{t}\,T\rfloor.$
where $N_{t}$ is the integer count of timesteps to mask for task $t$, computed as the desired masking proportion $p_{t}$ times the window length $T$ and then rounded down.
A.3 Span‑Based Poisson Masking (channel-aligned)
With $\lambda=7$:
  1. 1.
    Draw span length $L\sim\mathrm{Poisson}(\lambda)$, truncate to $1\leq L<T$.
  2. 2.
    Draw start index $s\sim\mathrm{U}\{0,\dots,T-L\}$.
  3. 3.
    Add span $S=\{s,\dots,s+L-1\}$ to the appropriate mask set(s) for the task.
Channel alignment means
$\mathcal{M}_{E}=\{(t,c)\mid t\in S,\;c\in\mathcal{C}\},\quad\mathcal{C}=\{1,\dots,C\}.$
Repeat until $|\mathcal{M}_{\{\cdot\}}|\geq N_{t}$.
A.4 End‑of‑Window Masking
If “End” is chosen,
$\mathcal{M}_{E}=\{(t,c)\mid t>T-N_{t},\;c\in\mathcal{C}\},\qquad\mathcal{M}_{A}=\{t\mid t>T-N_{t}\}.$
A.5 Targeted Transition Masking (task 0)
With buffer $B=50$ and first transition $\tau$ such that $B\leq\tau\leq T-B$:
$\Pr(\text{Targeted})$ $=0.5,$ $\mathcal{M}_{A}$ $=\{\,t\mid|t-\tau|<B\,\},\;\mathcal{M}_{E}=\varnothing;$
$\Pr(\text{End, }p_{t}=1)$ $=0.5.$
A.6 Self‑Supervised EMG Reconstruction (task 3)
A.6.1 Mask generation
Always Poisson, channel-aligned, with $p_{3}\!\in\![0.20,0.65]$. Attention flow to action is restricted via an attention mask.
A.6.2 Objective
$\mathcal{L}_{\mathrm{SS}}=\frac{1}{|\mathcal{M}_{E}|}\!\sum_{(t,c)\in\mathcal{M}_{E}}\bigl\|f_{E}(\tilde{x})_{t,c}-x_{t,c}\bigr\|^{2}_{2}$
A.7 Mask Tokens & Model Interface
  • Action mask token: $\langle\!\text{MASK}\!\rangle=K=3$ (in the 3-class case).
  • EMG mask token: numerical values remain unchanged; the boolean mask instructs the model to swap the corresponding embeddings with a learned vector $\mathbf{m}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension.
The model therefore consumes
$\tilde{x}_{t,c}=\begin{cases}\mathbf{m},&(t,c)\in\mathcal{M}_{E},\\ x_{t,c},&\text{otherwise},\end{cases}\qquad\tilde{y}_{t}=\begin{cases}\langle\!\text{MASK}\!\rangle,&t\in\mathcal{M}_{A},\\ y_{t},&\text{otherwise}.\end{cases}$
A.8 Composite Loss
$\mathcal{L}=\!\!\sum_{(t,c)\in\mathcal{M}_{E}}\frac{\|f_{E}(\tilde{x},\tilde{y})_{t,c}-x_{t,c}\|^{2}_{2}}{|\mathcal{M}_{E}|}\;+\;\sum_{t\in\mathcal{M}_{A}}\frac{\mathrm{CE}\!\bigl(f_{A}(\tilde{x},\tilde{y})_{t},\,y_{t}\bigr)}{|\mathcal{M}_{A}|}.$
For task 3 the CE term is zero.
Appendix B Online Inference
@@
Figure 4 illustrates our method as used for online inference. In order to label the intent at time $t$, we take all the overlapping windows of size equal to 600 timesteps = $3s$ and whose final sample falls inside the look-ahead horizon $t,t+\ell]$ (yellow region, $\ell=50$ timesteps = 250 ms), and feed them through our model. Our model then produces one logit for every timestep of each of these windows. We then average all the logits for timesteps falling within the lookahead horizon, and apply an $\arg\max$ to the result in order to obtain a single class label for time step $t$. To avoid jitter, the resulting label from time $t$ is then maintained for a hold period of $s=20$ timesteps or 100 ms, (blue region), after which the online inference process is performed again for time $t+100ms$.
@@
We note that our approach inherently introduces two types of latency in the response. First, the label for time $t$ is issued after ingesting data up to $t+250ms$, thus introducing up to 250ms delay in the response to a change in intent. However, this delay is smaller than the inherent delay in human reaction time to visual or verbal cues (which we discuss further in Section G). Second, the hold period means that our model will not react to intent changes happening at a rate faster than 10 Hz. However, we assume it to be highly unlikely that a human subject will modulate their hand pose or intent at such a rate.
Figure 4: Online inference with look-ahead horizon and reduced inference frequency. Five representative overlapping sliding windows are depicted, ending at time steps $t,t{+}10,\dots,t{+}50$ respectively. The small grey vertical bars within each window indicate the per-timestep logits that make up the average. All models use window size 600, with look-ahead 50 timesteps (0.25s) and a label hold time of 20 timesteps (0.1s)
Appendix C Dataset Description
@@
Figure 5 compares representative predictions from two models: one trained only on the public datasets and another pretrained on the public datasets and fine-tuned on our data. We evaluate both models on two tasks from our dataset: unsupported static hand gestures and functional grasp of a soup can. As shown, while both models initially identify the correct intent, the pretrain-only model quickly reverts to a “rest” state. In contrast, the fine-tuned model maintains the correct gesture label throughout the task, demonstrating improved temporal consistency and real-world applicability.
Appendix D Model Hyperparameters and Training Configurations
@@
Appendix E Baseline Implementations
@@
Betthauser et al. [17] proposes ED-TCN, an encoder–decoder temporal convolutional network that relies solely on 1D convolutions. To establish a fair basis of comparison, all baseline models use 600 window size, except ED-TCN. To align with the original paper’s hyperparameters, we increase the ED-TCN context length from 600 to 625 samples. The model processes these 625-sample “outer windows” with a stride of 30. Within each outer window, we then slide a 150-sample “inner window” (stride 25), producing 20 sub-windows (as opposed to 19 if the length remained 600). For each sub-window, we compute the mean absolute value (MAV) across all eight EMG channels, resulting in a 20×8 tensor that serves as input to ED-TCN. We maintain the original two-layer encoder structure, where each layer consists of a one-dimensional convolution, ReLU activation, and 2× max pooling. The decoder mirrors this architecture by applying 2× upsampling and the same convolution+ReLU sequence, concluding with a 1×1 convolution that outputs class probabilities for each time step. To accommodate the 625-sample windows, we scale down the convolution kernel size proportionally, using a kernel size of 9. This ensures the final temporal length is exactly 20 after pooling and upsampling. During training, we compute cross-entropy loss at each sub-window and backpropagate through every time step (where each time step corresponds to the MAV values of a sub-window), and use the AdamW optimizer (learning rate = 1e‑4) to mitigate overfitting. This slight adjustment—using 625-sample windows instead of 600 and a kernel size of 9 instead of 25—aligns the number of sub-windows with ED-TCN’s down/up‑sampling factors, ensuring a one-to-one mapping between sub-window features and labels without additional padding or cropping.
@@
Appendix F Context Window Size Ablation
Window Size. Overall window size controls the amount of context that the model sees when making a prediction. We vary the window size from 100 timesteps (0.5s) to 600 timesteps (3s) to assess how temporal context affects intent detection. Table 5 shows results of our model using three different window sizes on the EMG-EPN-612 6‑class dataset, with the evaluation configured to use lookahead 0 so that we can directly compare raw performance without any smoothing effects. Smaller windows reduce predictive stability, likely because the model lacks sufficient information to reliably distinguish transitory muscle signals. Consequently, we select a 600-timestep window size in our final model.
@@ Appendix F Context Window Size Ablation
Appendix G Reaction Time Discussion
@@
Ground truth intent labels are typically provided by a human experimenter, who cues the subject to execute a given command, while simultaneously marking the new command as ground truth intent. However, human reaction times introduce a baseline delay in muscle activation relative to visually or verbally cued labels. Simple reaction times to visual stimuli commonly range from about 200–250 ms, though they can be shorter in younger adults or under optimal conditions, and can approach 300 ms if the individual is older, fatigued, or otherwise slower to respond [56, 57]. Auditory cues tend to elicit slightly faster reaction times (around 150–200 ms) because neural processing for sound is somewhat quicker than for visual signals [58, 57]. Furthermore, participants may anticipate cues, activating muscles before or after the ground-truth label shift, which also misaligns the recorded EMG signal with the labeled transition. Manually labeling EMG datasets offline is imprecise and subjective, particularly when transitioning between non-rest states. Consequently, training and validation data can contain a baseline offset of 200–300 ms between the true transition and the labeled one. To account for this, we report results on our transition accuracy metric using a default reaction buffer size of 100 timesteps =500ms=500ms on each side of the ground truth labeled transition, comparable to the natural variability in reaction times and labeling inconsistencies. We also note that the same buffer size is used for all compared models, both ours and baselines.
@@
We also note that our model adds only negligible computational delay, beyond the lookahead window discussed in Sec. B, as it requires no additional data post processing. On an Nvidia RTX 4060 laptop GPU, the forward inference time is approximately 3 ms, which we consider negligible compared to the physiological and labeling-induced delays discussed above.
@@
Appendix H Random Seed Experiments
In Table 7, we show raw accuracy and transition accuracy results averaged over five different random seeds, along with standard deviation, for both our method and all baselines, evaluated on the EPN 6‑class dataset, with no lookahead. We notice that all models appear largely insensitive to the choice of random seed.